OpenStudy (anonymous):

Combustion analysis of an unknown compound provides the following data: 73.5 grams C, 4.20 grams H, and 72.3 grams Cl. What is the percent composition of each element in this compound? a. 73.5% C, 4.20% H, 72.3% Cl b.48.0% C, 3.00% H, 50.0% Cl c. 17.5% C, 1.00% H, 17.2% Cl d. 49.0% C, 2.80% H, 48.2% Cl

OpenStudy (anonymous):

ok firs divide each mass with matching atomic mass: 73,5 g / 12,011 =6,12 4,2 g / 1,008 = 4,16 72,3 g / 35,45 = 2,04 so now you know it is dichlor benzene:  now by doing same steps as in last task calculate percentages.